68 coins, all of which are 20 c or 50c pieces, amount to $25.30. find how many there are of each coin. Please tell me working out too.
Answer by stanbon(75887)

68 coins, all of which are 20 c or 50c pieces, amount to $25.30. find how many there are of each coin.
--------------------
Quantity Eq::: t + f = 68 coins
Value Eq:::: 20t + 50f = 2530 cents
------------
Multiply 1st thru by 20:
20t + 20f = 20*68
20t + 50f = 2530
-----------------------
Subtract and solve for "f":
30f = 1170
---
f = 39 (# of 50c coins)
------
t = 68-39 = 29 (# of 20c coins)
====================================
Cheers,
Stan H.
==============